OpenStudy (anonymous):

Solve the quadratic equation by the square root property. 5(x - 9)2 = 30

OpenStudy (anonymous):

Suppose to be 5(x-9)^2=30

OpenStudy (anonymous):

\[5 (x-9 ) ^2=30\\ (x-9 ) ^2=\frac{ 30 }{ 5 }\\ (x-9 ) ^2=6\\ x-9 =\pm \sqrt{6}\\ x= 9 \pm \sqrt{6}\]